How do you write the following question in point-slope form? Goes through the points (4,2) parellel to y=-3/4x-5
Answer by stanbon(75887)

How do you write the following question in point-slope form? Goes through the points (4,2) parellel to y=-3/4x-5
---
Parallel to that line means the slope must be -3/4
--------------------------------------------------
Eq. Form: y = mx + b
2 = (-3/4)*4 + b
b = 2+3
b = 5
-----
Equation:
y = (-3/4)x + 5
=================
Cheers,
Stan H.
=================